# Spectral Convolutional Conditional Neural Processes

### v4

@@ Appendix A Discretization Sensitivity of the DFT
To illustrate the sensitivity of the discrete Fourier transform (DFT) to grid shape—and how changes in discretization can disturb predictive performance—we consider a simple one-dimensional example (although the analysis extends to higher dimensions). Let $0<\Delta\ll 1$ denote a fixed discretization resolution, interpreted as the spacing between successive points. Define the grid $\mathcal{G}_{1}=\left(x_{0},\dots,x_{m_{1}-1}\right)$ as a uniform discretization of the interval $[0,1]$ with $m_{1}=\lfloor\frac{1}{\Delta}\rfloor$ and ordered such that $x_{k-1}<x_{k}$ for all $k\in\{1,\dots,m_{1}-1\}$. Let the sampled values of $\varphi_{e}[\mathcal{D}_{c}]$ at $\mathcal{G}_{1}$ be denoted by $\left(\varphi_{e}\left[\mathcal{D}_{c}\right](x)\right)_{x\in\mathcal{G}_{1}}=\left(\varphi_{e}[\mathcal{D}_{c}](x_{0}),\;\dots\;,\varphi_{e}[\mathcal{D}_{c}](x_{m_{1}-1})\right)$. The DFT of $\varphi_{e}[\mathcal{D}_{c}]$ over this grid is given by:
@@ B.1.1 Model Architectures
This section details the architectures of all CNP variants used in our experiments. Each model outputs a Gaussian predictive distribution parameterized by a mean and a pre-softplus [dugas2000incorporating] scale. The pre-softplus value is passed through a softplus transformation, and a minimum noise term of $10^{-6}$ is then added to the resulting scale. The Gaussian factorizes across query points—as is standard for CNPs (see Section 2.2)—and, for multidimensional outputs, across output dimensions as well. Thus, predictions assume independence both across query locations and across components of each output vector. Unless otherwise noted, all nonlinearities use ReLU activations [nair2010rectified].
@@
The CNP encodes each context pair $(x_{c,k},y_{c,k})\in\mathcal{D}_{c}$ using separate input and output pathways. The input $x_{c,k}$ and output $y_{c,k}$ are first passed through distinct MLPs with two hidden layers each of dimension 256, producing 256-dimensional representations $\varepsilon^{(x)}_{c,k}$ and $\varepsilon^{(y)}_{c,k}$. Their concatenation $\varepsilon_{c,k}=[\varepsilon^{(x)}_{c,k},\varepsilon^{(y)}_{c,k}$] is then processed by a deeper MLP with six hidden layers of width 256 and a final 256-dimensional output layer, yielding an embedding for each context pair. Context embeddings $\varepsilon_{c,k}$ are averaged to form a single aggregated representation $\varepsilon_{c}$ of dimension 256. For prediction, each query input $x_{q,k}$ is concatenated with the context representation $\varepsilon_{c}$ and passed through a decoder MLP consisting of six hidden layers of dimension 256. The decoder’s final output layer has dimensionality $2d_{{\mathbb{Y}}}$, parameterizing the mean and the log-scale parameters of the Gaussian predictive distribution.
@@
Our AttCNP implementation follows the deterministic architecture introduced by kim2019attentive. The initial encoding stage mirrors that of the CNP: each context pair $(x_{c,k},y_{c,k})\in\mathcal{D}_{c}$ is processed by two separate MLPs (each with two hidden layers of width 256), producing embeddings $\varepsilon^{(x)}_{c,k}$ and $\varepsilon^{(y)}_{c,k}$. These are concatenated to form $\varepsilon_{c,k}=[\varepsilon^{(x)}_{c,k},\varepsilon^{(y)}_{c,k}$], which is subsequently passed through an additional two-layer MLP with hidden width 256. Departing from the CNP, the AttCNP applies self-attention to the set of context embeddings. We use two layers of multi-head self-attention, each with 8 heads (head dimension 32), a feedforward subnetwork with hidden width 256, residual connections, and pre-layer normalization for both the attention and feedforward blocks. The input and output dimensionality of each attention layer is fixed at 256. Query representations $\varepsilon_{q,k}$ are computed using a single layer of multi-head cross-attention between the (attended) context embeddings and the query embeddings, with the same configuration as above (8 heads of dimension 32, feedforward width 256, residual connections, and pre-normalization). The decoder is an MLP with six hidden layers of width 256. Its final layer outputs $2d_{{\mathbb{Y}}}$ units, corresponding to the mean and log-scale parameters of a Gaussian predictive distribution.
@@
where the zero-vector matches the shape of $y_{c,k}$ and the final “0” indicates the absence of an observation [nguyen2022transformer, ashman2024translation]. All tokens are then passed through a shared two-layer MLP (width 256), producing initial embeddings $\varepsilon_{c,k}$ and $\varepsilon_{q,k}$.
@@
Our model uses six transformer layers with eight attention heads (head dimension 32), each followed by a feedforward subnetwork of width 256. We use the standard pre-norm architecture, applying layer normalization before both the attention and feedforward modules. All attention operations share a 256-dimensional embedding dimension. At each layer, we first apply context self-attention and then query–context cross-attention. Notably, the same multi-head attention parameters are used for both operations; the distinction arises only from which embeddings serve as queries, keys, and values (yielding self-attention when they coincide, and cross-attention otherwise). This is not a modeling requirement—separate parameter sets could be used, as demonstrated in Appendix B.4.1. The final-layer query embeddings are passed to a decoder MLP with two hidden layers (width 256), whose output parameterizes a Gaussian predictive distribution via $2d_{{\mathbb{Y}}}$ units corresponding to the mean and log-scale.
Translation Equivariant Transformer Neural Process (TETNP).
The token construction for TETNP closely mirrors that of the TNP, with the key difference that token representations omit the input locations in order to preserve translation equivariance. For each context pair $(x_{c,k},y_{c,k})\in\mathcal{D}_{c}$, we therefore construct the token
@@
and for each query location $x_{q,k}$ we use
@@
All tokens are processed by a shared two-layer MLP (width 256), yielding initial embeddings $\varepsilon_{c,k}$ and $\varepsilon_{q,k}$. TETNP replaces the standard multi-head attention mechanism with the translation-equivariant variant introduced by ashman2024translation. To compute attention scores, the pairwise scaled dot-product between token embeddings for each head is concatenated with the corresponding pairwise differences in token locations. This concatenated feature is passed through an MLP with two hidden layers (width 256) and an output dimension equal to the number of attention heads, producing translation-equivariant attention logits. Except for this modified attention module, the remaining architecture and computation flow follow the implementation of the TNP.
@@
For ConvCNP, we begin by determining, for each input dimension, the minimum and maximum coordinates observed across both the context and query sets. These extrema are expanded by a small margin of 0.1 in each dimension. The interval between the expanded minima and maxima is then uniformly discretized at a resolution of 64 points per unit. When necessary—for example, to satisfy the CNN’s minimum grid-size requirements—the discretization range is further enlarged while maintaining the same resolution. The resulting one-dimensional grids are combined via a Cartesian product to yield a uniform $d_{{\mathbb{X}}}$-dimensional grid $\mathcal{G}$.
The functional embedding in equation 3 is evaluated on this grid. The Gaussian kernels used in this embedding are initialized with per-dimension length scales set to twice the grid resolution, i.e., $2/64$. For multi-dimensional inputs and outputs, the model uses separate length scales for each dimension. The embedding values are further normalized using a density channel, defined as
@@
The functional and density channels are concatenated to form the representation
$\left(\mathrm{Density}(x_{g}),\ \frac{\sum_{(x_{c},\,y_{c})\in\mathcal{D}_{c}}\phi(y_{c})\,\psi_{e}(x_{g}-x_{c})}{\sum_{(x_{c},\,y_{c})\in\mathcal{D}_{c}}\psi_{e}(x_{g}-x_{c})}\right)_{x_{g}\in\mathcal{G}}.$ (5)
Each grid location is processed by an MLP with two hidden layers of width 128. The resulting features are passed to a CNN based on a U-Net architecture [ronneberger2015u], consisting of six residual convolutional blocks (kernel size 11, stride 2, 128 channels) in the encoder, followed by a symmetric sequence of six transposed convolutional blocks in the decoder, with skip connections following the design of bruinsma2023autoregressive. Because the U-Net downsamples the spatial resolution by a factor of 64, we ensure that the constructed grid size is divisible by 64. For the one-dimensional benchmarks considered here, we therefore enlarge the discretization interval when necessary so that the CNN’s output aligns exactly with the input grid size.
To obtain predictive distribution parameters at query locations—which may not lie on the grid $\mathcal{G}$— an interpolation scheme analogous to equation 5 is applied, except that the weighted sum over the feature-map values is not normalized. For this interpolation, we employ a separate Gaussian kernel with learnable length-scale parameters, distinct from the kernel used in the functional embedding. The resulting query-specific embeddings are then processed by a decoder MLP with two hidden layers (width 128), whose output consists of $2d_{{\mathbb{Y}}}$ units parameterizing the mean and log-scale of a Gaussian predictive distribution.
@@
Table 5 summarizes the number of learnable parameters for all models evaluated in our 1D synthetic regression experiments.
Table 5: Learnable parameter counts for all models evaluated in the 1D synthetic regression experiments.
@@
Table 6: Forward-pass runtime (in seconds) for a batch of 16 tasks during training and validation across all models in the 1D synthetic regression setting.
@@
B.1.2 Training and Data Description
@@
Sawtooth $f(x)=2\Big(\big(\xi(u^{\top}x)-c\big)\bmod 1\Big)+\sigma_{0}(x),\,\,\sigma_{0}(x)\sim\mathcal{N}(0,\sigma_{0}^{2})$
Square Wave $f(x)=2\,\mathds{1}_{\Big\{\big((\xi x-c)\bmod 1\big)<D\Big\}}-1+\sigma_{0}(x),\,\,\sigma_{0}(x)\sim\mathcal{N}(0,\sigma_{0}^{2})$
@@
For validation, we use a fixed set of 4,096 tasks, organized into 256 batches of 16 tasks. In these tasks, the number of query points is fixed at 256, while the number of context points is sampled using the same procedure as during training. Testing follows the same configuration as validation, except that we evaluate on 16,000 test tasks. Unlike the dynamically generated training tasks, the validation and test sets remain fixed across all experiments.
@@ B.2.1 Model Architectures
The model architectures and parameter counts closely follow those described in B.1.1. The only modification concerns the discretization of the functional embedding in ConvCNP and SConvCNP. Specifically, we first expand the range defined by the minima and maxima of all context and query coordinates by a margin of 0.5 in each dimension. The resulting interval is then uniformly discretized at a resolution of 48 points per unit length.
B.2.2 Training and Data Description
@@
$dU_{t}$ $=\alpha U_{t}\,dt-\beta U_{t}V_{t}\,dt+\sigma U_{t}^{\nu}\,dB^{(1)}_{t}$,
$dV_{t}$ $=-\gamma V_{t}\,dt+\delta U_{t}V_{t}\,dt+\sigma V_{t}^{\nu}\,dB^{(2)}_{t}$,
where $B^{(1)}_{t}$ and $B^{(2)}_{t}$ are independent Brownian motions. The deterministic drift recovers classical predator–prey behavior: prey grow exponentially at rate $\alpha$ in the absence of predators, while predators decline at rate $\gamma$ without prey. The interaction terms $\beta U_{t}V_{t}$ and $\delta U_{t}V_{t}$ model predation and predator reproduction. Stochasticity enters through the multiplicative noise terms $\sigma U_{t}^{\nu}$ and $\sigma V_{t}^{\nu}$, where $\sigma$ controls noise intensity and $\nu$ determines how fluctuations scale with population size. Setting $\nu=1$ yields noise proportional to population level, while $\nu<1$ and $\nu>1$ induce sublinear and superlinear scaling, respectively.
@@
For validation, we use a fixed set of 4,096 tasks, arranged into 256 batches of 16 tasks. In these tasks, the number of query points is fixed at 256, while the number of context points is sampled using the same procedure as in training. Testing follows the same protocol as validation, except that we evaluate on 16,000 test tasks.
To construct a task from a simulated trajectory, input locations $t$ are sampled uniformly and independently from the recorded time points in $[0,100)$ and paired with the corresponding values $(U_{t},V_{t})$. We rescale time by a factor of 0.1, mapping the 100-year interval to $[0,10)$, and rescale population sizes by multiplying the values by 0.01. Unlike the dynamically generated training tasks, the randomization used to construct the validation and test tasks is fixed across all experiments.

### v5

@@ Appendix A Discretization Sensitivity of the DFT
To illustrate the sensitivity of the discrete Fourier transform (DFT) to grid shape—and how changes in discretization can disturb predictive performance—we consider a simple one-dimensional example. Let $0<\Delta\ll 1$ denote a fixed discretization resolution, interpreted as the spacing between successive points. Define the grid $\mathcal{G}_{1}=\left(x_{0},\dots,x_{m_{1}-1}\right)$ as a uniform discretization of the interval $[0,1]$ with $m_{1}=\lfloor\frac{1}{\Delta}\rfloor$ and ordered such that $x_{k-1}<x_{k}$ for all $k\in\{1,\dots,m_{1}-1\}$. Let the sampled values of $\varphi_{e}[\mathcal{D}_{c}]$ at $\mathcal{G}_{1}$ be denoted by $\left(\varphi_{e}\left[\mathcal{D}_{c}\right](x)\right)_{x\in\mathcal{G}_{1}}=\left(\varphi_{e}[\mathcal{D}_{c}](x_{0}),\;\dots\;,\varphi_{e}[\mathcal{D}_{c}](x_{m_{1}-1})\right)$. The DFT of $\varphi_{e}[\mathcal{D}_{c}]$ over this grid is given by:
@@ B.1.1 Model Architectures
This section details the architectures of all CNPs variants used in our experiments. Each model outputs a Gaussian predictive distribution parameterized by a mean and a pre-softplus [dugas2000incorporating] scale. The pre-softplus value is passed through a softplus transformation, and a minimum noise term of $10^{-6}$ is then added to the resulting scale. The Gaussian factorizes across query points—as is standard for CNPs (see Section 2.2)—and, for multidimensional outputs, across output dimensions as well. Thus, predictions assume independence both across query locations and across components of each output vector. Unless otherwise noted, all nonlinearities use ReLU activations [nair2010rectified].
@@
The CNP encodes each context pair $(x_{c,k},y_{c,k})\in\mathcal{D}_{c}$ using separate input and output pathways. The input $x_{c,k}$ and output $y_{c,k}$ are first passed through distinct MLPs with two hidden layers each of dimension 256, producing 256-dimensional representations $\varepsilon^{(x)}_{c,k}$ and $\varepsilon^{(y)}_{c,k}$. Their concatenation $\varepsilon_{c,k}=[\varepsilon^{(x)}_{c,k},\varepsilon^{(y)}_{c,k}$] is then processed by a deeper MLP with six hidden layers of width 256 and a final 256-dimensional output layer, yielding an embedding for each context pair. Context embeddings $\varepsilon_{c,k}$ are averaged to form a single aggregated representation $\varepsilon_{c}$ of dimension 256. For prediction, each query input $x_{q,k}$ is concatenated with the context representation $\varepsilon_{c}$ and passed through a decoder MLP consisting of six hidden layers of dimension 256. The decoder’s final output layer has dimensionality $2d_{y}$, parameterizing the mean and the log-scale parameters of the Gaussian predictive distribution.
@@
The AttCNP implementation follows the deterministic architecture introduced by kim2019attentive. The initial encoding stage mirrors that of the CNP: each context pair $(x_{c,k},y_{c,k})\in\mathcal{D}_{c}$ is processed by two separate MLPs (each with two hidden layers of width 256), producing embeddings $\varepsilon^{(x)}_{c,k}$ and $\varepsilon^{(y)}_{c,k}$. These are concatenated to form $\varepsilon_{c,k}=[\varepsilon^{(x)}_{c,k},\varepsilon^{(y)}_{c,k}$], which is subsequently passed through an additional two-layer MLP with hidden width 256. Departing from the CNP, the AttCNP applies self-attention to the set of context embeddings. We use two layers of multi-head self-attention, each with 8 heads (head dimension 32), a feedforward subnetwork with hidden width 256, residual connections, and pre-layer normalization for both the attention and feedforward blocks. The input and output dimensionality of each attention layer is fixed at 256. Query representations $\varepsilon_{q,k}$ are computed using a single layer of multi-head cross-attention between the (attended) context embeddings and the query embeddings, with the same configuration as above (8 heads of dimension 32, feedforward width 256, residual connections, and pre-normalization). The decoder is an MLP with six hidden layers of width 256. Its final layer outputs $2d_{y}$ units, corresponding to the mean and log-scale parameters of a Gaussian predictive distribution.
@@
where the dummy zero-vector matches the shape of $y_{c,k}$ and the final “0” indicates the absence of an observation [nguyen2022transformer, ashman2024translation]. All tokens are then passed through a shared two-layer MLP (width 256), producing initial embeddings $\varepsilon_{c,k}$ and $\varepsilon_{q,k}$.
@@
Our model uses six transformer layers with eight attention heads (head dimension 32), each followed by a feedforward subnetwork of width 256. We use the standard pre-norm architecture, applying layer normalization before both the attention and feedforward modules. All attention operations share a 256-dimensional embedding dimension. At each layer, we first apply context self-attention and then query–context cross-attention. Notably, the same multi-head attention parameters are used for both operations; the distinction arises only from which embeddings serve as queries, keys, and values (yielding self-attention when they coincide, and cross-attention otherwise). This is not a modeling requirement—separate parameter sets could be used, as demonstrated in Appendix B.4.1. The final-layer query embeddings are passed to a decoder MLP with two hidden layers (width 256), whose output parameterizes a Gaussian predictive distribution via $2d_{y}$ units corresponding to the mean and log-scale.
Translation-Equivariant Transformer Neural Process (TETNP).
The token construction in the TETNP closely follows that of the TNP, with a crucial modification to ensure translation equivariance: input locations are excluded from the token representations. Specifically, for each context pair $(x_{c,k},y_{c,k})\in\mathcal{D}_{c}$, we construct the token
@@
while for each query location $x_{q,k}$ we use the token
@@
All tokens are passed through a shared two-layer MLP with width 256 to produce initial embeddings $\varepsilon_{c,k}$ and $\varepsilon_{q,k}$.
The TETNP replaces the standard multi-head attention mechanism with the translation-equivariant attention proposed by ashman2024translation. For each attention head, we first compute the matrix of pairwise scaled dot products between token embeddings, analogous to conventional attention but independent of absolute input locations. In parallel, we compute the matrix of pairwise differences between token locations. We then concatenate the scaled dot-product scores from all heads with the corresponding pairwise location differences. This augmented pairwise similarity representation is processed by an MLP—acting as an implicit kernel—with two hidden layers of width 256 and an output dimension equal to the number of attention heads, yielding the final translation-equivariant attention logits.
Apart from this modified attention module, the remainder of the architecture and computational pipeline follows that of the TNP.
@@
For the ConvCNP, we begin by determining, for each input dimension, the minimum and maximum coordinates observed across both the context and query sets. These extrema are expanded by a small margin of 0.1 in each dimension. The interval between the expanded minima and maxima is then uniformly discretized at a resolution of 64 points per unit. When necessary—for example, to satisfy the CNN’s minimum grid-size requirements—the discretization range is further enlarged while maintaining the same resolution. The resulting one-dimensional grids are combined via a Cartesian product to yield a uniform grid $\mathcal{G}$.
The functional embedding in equation 3 is evaluated on this grid. The Gaussian kernels used in this embedding are initialized with per-dimension length scales set to twice the grid resolution, i.e., $2/64$. For multi-dimensional inputs and outputs, the model uses separate length scales for each dimension. Following bruinsma2023autoregressive, the embedding is further normalized using a density channel, defined as
@@
The functional and density channels are concatenated to form the final grid representation
$\left(\mathrm{Density}(x_{g}),\ \frac{\sum_{(x_{c},\,y_{c})\in\mathcal{D}_{c}}\phi(y_{c})\,\psi_{e}(x_{g}-x_{c})}{\mathrm{Density}(x_{g})}\right)_{x_{g}\in\mathcal{G}}.$ (5)
Each grid point is processed independently by an MLP with two hidden layers of width 128. The resulting features are passed to a CNN based on a U-Net architecture [ronneberger2015u], consisting of six residual convolutional blocks (kernel size 11, stride 2, 128 channels) in the encoder, followed by a symmetric sequence of six transposed convolutional blocks in the decoder, with skip connections following the design of bruinsma2023autoregressive. Since the U-Net downsamples the spatial resolution by a factor of 64, we ensure that the constructed grid size is divisible by 64. For the one-dimensional benchmarks considered in this work, this is achieved by symmetrically enlarging the discretization interval when necessary so that the CNN output aligns exactly with the input grid.
To obtain predictive distribution parameters at query locations—which may not lie on the grid $\mathcal{G}$— an interpolation scheme analogous to equation 5 is applied, except that the weighted sum over the feature-map values is not normalized. For this interpolation, we employ a separate Gaussian kernel with learnable length-scale parameters, distinct from the kernel used in the functional embedding. The resulting query-specific embeddings are then processed by a decoder MLP with two hidden layers (width 128), whose output consists of $2d_{y}$ units parameterizing the mean and log-scale of a Gaussian predictive distribution.
@@
Table 5 summarizes the number of learnable parameters for all models used in our 1D synthetic regression experiments.
Table 5: Learnable parameter counts for all models used in the 1D synthetic regression experiments.
@@
Table 6: Average forward-pass runtime (in seconds) for a batch of 16 tasks during training and validation across all models in the 1D synthetic regression setting.
@@
B.1.2 Data and Experimental Setup
@@
Sawtooth $f(x)=2\left(\left(\xi\left(u^{\top}x-c\right)\right)\bmod 1\right)+\sigma_{0}(x),\,\,\sigma_{0}(x)\sim\mathcal{N}(0,\sigma_{0}^{2})$
Square Wave $f(x)=2\,\mathds{1}_{\left\{\left(\left(\xi x-c\right)\bmod 1\right)<D\right\}}-1+\sigma_{0}(x),\,\,\sigma_{0}(x)\sim\mathcal{N}(0,\sigma_{0}^{2})$
@@
For validation, we use a fixed set of 4,096 tasks, organized into 256 batches of 16 tasks. In these tasks, the number of query points is fixed at 256, while the number of context points is sampled using the same procedure as during training. Testing follows the same configuration as validation, except that we evaluate on 16,000 test tasks. Unlike the dynamically generated training tasks, the validation and test sets remain fixed across all experiments and runs.
@@ B.2.1 Model Architectures
The model architectures and parameter counts closely follow those described in B.1.1. The only modification concerns the discretization of the functional embedding in the ConvCNP and the SConvCNP. Specifically, we first expand the range defined by the minima and maxima of all context and query coordinates by a margin of 0.5 in each dimension. The resulting interval is then uniformly discretized at a resolution of 48 points per unit length.
B.2.2 Data and Experimental Setup
@@
$dU_{t}$ $=\alpha U_{t}\,dt-\beta U_{t}V_{t}\,dt+\sigma U^{\nu}_{t}\,dB^{(1)}_{t}$,
$dV_{t}$ $=-\gamma U_{t}\,dt+\delta U_{t}V_{t}\,dt+\sigma V^{\nu}_{t}\,dB^{(2)}_{t}$,
where $B^{(1)}_{t}$ and $B^{(2)}_{t}$ are independent Brownian motions. The deterministic drift recovers classical predator–prey behavior: prey grow exponentially at rate $\alpha$ in the absence of predators, while predators decline at rate $\gamma$ without prey. The interaction terms $\beta U_{t}V_{t}$ and $\delta U_{t}V_{t}$ model predation and predator reproduction. Stochasticity enters through the multiplicative noise terms $\sigma U^{\nu}_{t}$ and $\sigma V^{\nu}_{t}$, where $\sigma$ controls noise intensity and $\nu$ determines how fluctuations scale with population size. Setting $\nu=1$ yields noise proportional to population level, while $\nu<1$ and $\nu>1$ induce sublinear and superlinear scaling, respectively.
@@
To construct a task from a simulated trajectory, input locations $t$ are sampled uniformly and independently from the recorded time points in $[0,100)$ and paired with the corresponding values $(U_{t},V_{t})$. We rescale time by a factor of 0.1, mapping the 100-year interval to $[0,10)$, and rescale population sizes by multiplying the values by 0.01.
@@
For validation, we use a fixed set of 4,096 tasks, arranged into 256 batches of 16 tasks. In these tasks, the number of query points is fixed at 256, while the number of context points is sampled using the same procedure as in training. Testing follows the same protocol as validation, except that we evaluate on 16,000 test tasks. Unlike the dynamically generated training tasks, the randomization used to construct the validation and test tasks is fixed across all experiments and runs.